On January 1, 2019, Betty DeRose, Inc. purchased a piece of machinery for $320,000.
The machine had an estimated useful life of 8 years and a salvage value of $5,000.

Assume Betty DeRose depreciates its assets using the double declining balance method
of depreciation.

Calculate the amount of accumulated depreciation related to the machine that would
appear in Betty's December 31, 2022 balance sheet.
Answer #1

Accumulated depreciation at the end of December 31, 2022= $ 218,750

Working

Double declining Method
A Cost $     3,20,000.00
B Residual Value $           5,000.00
C=A - B Depreciable base $     3,15,000.00
D Life [in years] 8
E=C/D Annual SLM depreciation $        39,375.00
F=E/C SLM Rate 12.50%
G=F x 2 DDB Rate 25.00%
Year Beginning Book Value Depreciation rate Depreciation expense Ending Book Value Accumulated Depreciation
2019 $   3,20,000.00 25.00% $      80,000.00 $   2,40,000.00 $      80,000.00
2020 $   2,40,000.00 25.00% $      60,000.00 $   1,80,000.00 $   1,40,000.00
2021 $   1,80,000.00 25.00% $      45,000.00 $   1,35,000.00 $   1,85,000.00
2022 $   1,35,000.00 25.00% $      33,750.00 $   1,01,250.00 $   2,18,750.00